Cellulose could provide an abundant and cheap form of glucose for humans. Why is cellulose not a source of nutrients for humans? Cellulose digestion must occur in a basic environment, but the human gut is acidic. Humans, and most vertebrates, lack the enzyme cellulase. Humans cannot absorb the hydrolyzed form of cellulose Animals use starch and glycogen as an energy source, whereas plants use cellulose for energy. vertebrate enzymes are unable to hydrolyze (β 1-4) linked glucose monomers. □
Answer #1
Concepts and reason

The concept used here is based on the inability of humans to use cellulose as a source of nutrients.

Initially, the incorrect options are opted out regarding the digestion of cellulose. Then, the correct options are chosen which explains correctly the reason behind the inability of humans to use cellulose as a source of nutrients.

Fundamentals

Cellulose:

Cellulose is a polysaccharide made up of linear chains of monomers of β\beta - glucose which are linked together by (β14)\left( {\beta {\rm{ }}1 - 4} \right) glycosidic bonds. This results in a closely packed structure making it a vital component of the plant cell wall.

Cellulose is not soluble in weak acids or bases for its close-packed structure.

Product of hydrolysis of cellulose is glucose which is the main source of energy in humans.

Plants get energy from glucose which is produced by the process of photosynthesis.

Hence, the a,candda,{\rm{ }}c{\rm{ and }}d options are incorrect.

The enzymatic hydrolysis of cellulose can be done by using cellulase enzyme which is not present in humans and most vertebrates.

Glucose monomers with (β14)\left( {\beta {\rm{ }}1 - 4} \right) linkages form cellulose which cannot be hydrolyzed by the enzymes present in vertebrates.

Hence, the bandeb{\rm{ and }}e options are correct.

Ans:

Cellulose is not a source of nutrients for humans because humans, and most vertebrates, lack the enzyme cellulase and the vertebrate enzymes are unable to hydrolyze (β14)\left( {{\bf{\beta 1 - 4}}} \right) linked glucose monomers.